The area of a rectangle is 560 square inches. The length is 3 more than twice the width. Find the
length and the width.

Answer by TimothyLamb(4379)

---
Lw = 560
L = 2w + 3
---
Lw = 560
(2w + 3)w = 560
2ww + 3w - 560 = 0
---
the above quadratic equation is in standard form, with a=2, b=3, and c=-560
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 3 -560
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
w = 16
w = -17.5
---
negative width doesn't make sense for this problem, so use the positive root
---
answer:
w = 16 in
L = 35 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php